# Minimal Construction of Graphs with Maximum Robustness

### v1

@@ 6 Construction of $(\gamma,\gamma)$-MERGs
In the previous section, we have demonstrated the lower bounds on edge counts that any graph of $n$ nodes must satisfy in order to achieve $(\gamma,\gamma)$-robustness, where $\gamma=\lceil\frac{n}{2}\rceil$. Now, we present our second main result by constructing $(\gamma,\gamma)$-MERGs, or graphs that achieve $(\gamma,\gamma)$-robustness with a minimal set of edges:
Theorem 6.26.
@@
Proof 6.27.
Odd $n$: Because all $12\gamma-$ 1 nodes have $2(\gamma-1)$ neighbors, that means $\mathcal{G}$ forms a complete graph. Therefore by Lemma 5.18, $\mathcal{G}$ is a $(\gamma,\gamma)$-MERG when $n$ is odd.
Even $\mathbf{n}$: To prove that $\mathcal{G}$ is $(\gamma,\gamma)$-MERG for even values of $n$, we need to prove two things: (i) it is $(\gamma,\gamma)$-robust and (ii) its edge count $|\mathcal{E}|$ equals to the lower bounds specified in Lemma 5.22.
First, we prove its robustness. WLOG, we form disjoint non-empty subsets $\mathcal{S}_{1},\mathcal{S}_{2}\subset\mathcal{V}$ such that $|\mathcal{S}_{1}|\leq|\mathcal{S}_{2}|$, which implies $1\leq|\mathcal{S}_{1}|\leq\gamma$. By construction, there are at least $\gamma$ nodes in $\mathcal{G}$ that are connected to all other nodes (i.e., they has edges with $12\gamma-$ 1 nodes). There are two different cases:
@@
2) If $|\mathcal{S}_{1}|=|\mathcal{S}_{2}|=\gamma$, note that we have at least $\gamma$ nodes that have $12\gamma-$ 1 neighbors. Thus, however we form $\mathcal{S}_{1}$ and $\mathcal{S}_{2}$, there will be at least $\gamma=\frac{n}{2}$ nodes in $\mathcal{S}_{1}$ and $\mathcal{S}_{2}$ combined that have at least $2\gamma-1-(\gamma-1)=\gamma$ neighbors outside of $\mathcal{S}_{1}$ and $\mathcal{S}_{2}$, respectively. This satisfies the third condition given in Definition 8. Since $\mathcal{G}$ satisfies Definition 8 with $r=s=\gamma$ in all cases, $\mathcal{G}$ is $(\gamma,\gamma)$-robust.
Now we show that $\mathcal{E}$ is a minimal set. We have a total degree of $4\gamma(\gamma-1)+K$, where $K=\gamma+1$ if $\gamma$ is odd and or $K=\gamma$ if $\gamma$ is even. Thus, the total number of edges is $|\mathcal{E}|=\frac{4\gamma(\gamma-1)+K}{2}=\frac{K}{2}+2\gamma(\gamma-1)$, which equals the lower bound given in Lemma 5.22.
Theorem 6.26 addresses Problem 2 by constructing $(\gamma,\gamma)$-MERGs, graphs that achieve maximum $(r,s)$-robustness with a minimal set of edges. Similar to $\gamma$-MERGs, these graphs also have as many edges as specified by the lower bounds in Lemma 5.18 and 5.22. This shows that the lower bounds given in the lemmas are tight. In fact, removing any edge from $(\gamma,\gamma)$-MERGs immediately reduces the graphs’ robustness levels likewise.
Examples of $(\gamma,\gamma)$-MERGs as the results of Theorem 6.26 are visualized in Figure 4. The graph on the left has $n=9$ nodes and forms a complete graph to achieve $(5,5)$-robustness, as discussed in Lemma 5.18. On the contrary, the graph on the right has $n=10$ nodes to achieve $(5,5)$-robustness. Note, it contains an $8$-clique as the maximum clique (i.e., $\{0,3,4,5,6,7,8,9\}$), which validates Corollary 5.24.
Furthermore, we also highlight that the construction mechanisms given in Theorem 6.26 are necessary and sufficient in a sense that they encompass all the possible variations of $(\gamma,\gamma)$-MERGs.
@@
Let $\mathcal{G}=(\mathcal{V},\mathcal{E})$ be a graph of $n$ nodes, and let $\gamma=\lceil\frac{n}{2}\rceil$. Then, $\mathcal{G}$ is $(\gamma,\gamma)$-robust if and only if $\mathcal{G}$ contains a graph $\mathcal{G}^{\prime}=(\mathcal{V},\mathcal{E}^{\prime})$ from Theorem 6.26 as a spanning subgraph (i.e., $\mathcal{E}^{\prime}\subseteq\mathcal{E}$).
Proof 6.28.
@@
(necessary) Odd $n$: It is shown in Lemma 5.18 that complete graph is the only graph with $n$ nodes that is $(\gamma,s)$-robust where $s\geq\lfloor\frac{n}{2}\rfloor$. This indicates that $\mathcal{G}$ must be a complete graph, which means $\mathcal{E}^{\prime}=\mathcal{E}$, as $\mathcal{G}^{\prime}$ from Theorem 6.26 is also a complete graph.
@@
Even $n$: Here we prove the claim by contradiction. Assume by contradiction that $\mathcal{G}$ is $(\gamma,\gamma)$-robust and $\mathcal{E}^{\prime}\not\subseteq\mathcal{E}$. Let $\mathcal{N}_{i}$ and $\mathcal{N}_{i}^{\prime}$ be the neighbor sets of agent $i\in\mathcal{V}$ in $\mathcal{G}$ and $\mathcal{G}^{\prime}$, respectively. Then, we define =$\mathcal{M}=\{i\in\mathcal{V}\mid|\mathcal{N}_{i}|=2\gamma-1\}$, $\mathcal{P}=\{i\in\mathcal{V}\mid|\mathcal{N}_{i}|=2\gamma-2\}$, =$\mathcal{M}^{\prime}=\{i\in\mathcal{V}\mid|\mathcal{N}_{i}^{\prime}|=2\gamma-1\}$, and $\mathcal{P}^{\prime}=\{i\in\mathcal{V}\mid|\mathcal{N}_{i}^{\prime}|=2\gamma-2\}$.
@@
Note $\mathcal{G}^{\prime}$ has $|\mathcal{M}^{\prime}|=2\cdot\lceil\frac{\gamma}{2}\rceil$ nodes with degree of $12\gamma-$ 1 and $|\mathcal{P}^{\prime}|=2\gamma-|\mathcal{M}|$ nodes with degree of $22\gamma-$ 2. By Corollary 5.20, all nodes cannot have less than $22\gamma-$ 2 neighbors for $\mathcal{G}$ to be $(\gamma,\gamma)$-robust. Furthermore, they cannot have more than $12\gamma-$ 1 neighbors. What that means is that the only way for $\mathcal{E}^{\prime}\not\subseteq\mathcal{E}$ is to have $|\mathcal{M}|<|\mathcal{M}^{\prime}|$. Then, $|\mathcal{E}|<2\gamma(\gamma-1)+\lceil\frac{\gamma}{2}\rceil$. However, that is a contradiction, as for $\mathcal{G}$ to be $(\gamma,\gamma)$-robust, $|\mathcal{E}|\geq 2\gamma(\gamma-1)+\lceil\frac{\gamma}{2}\rceil$ by Lemma 5.22.
@@
(sufficiency) We know that $\mathcal{G}^{\prime}$ is a $(\gamma,\gamma)$-MERG, i.e., $(\gamma,\gamma)$-robust graph with a minimal set of edges. Since $(r,s)$-robustness is monotonic with respect to number of edges [4, Lemma 3], if $\mathcal{E}^{\prime}\subseteq\mathcal{E}$, $\mathcal{G}$ must be $(\gamma,\gamma)$-robust.
Proposition 1 strengthens Theorem 6.26, as it demonstrates that the construction mechanism in the theorem captures all variations of possible $(\gamma,\gamma)$-MERGs. We can also conclude that if a graph does not contain $(\gamma,\gamma)$-MERGs from Theorem 6.26 as a spanning subgraph, it is not $(\gamma,\gamma)$-robust. In other words, we can use Proposition 1 to quickly verify whether a given graph $\mathcal{G}$ is $(\gamma,\gamma)$-robust by checking its spanning subgraphs. Although detecting such subgraphs is, in general, an NP-complete problem, many efficient subgraph matching algorithms exist (e.g., [44, 45]). This result can be then used to upper-bound the value of $F$, the maximum number of malicious agents a given graph can tolerate, by upper-bounding its $(r,s)$-robustness without using algorithms like [20, 23], which scale poorly with the number of nodes. This is particularly valuable, especially because computing $(r,s)$-robustness is generally more challenging than computing $r$-robustness.
@@
Figure 6: Simulation results for the second set of simulations with $2$ Byzantine agents. We run consensus on $5$-MERGs with $n=9$ with and without an edge removal ((a) and (e)) and $n=10$ with and without an edge removal ((b) and (f)), and on $(5,5)$-MERGs with $n=9$ with and without an edge removal ((c) and (g)) and $n=10$ with and without an edge removal ((d) and (h)). The normal agents’ states are solid colored lines, while Byzantine agents’ states are shown as red dotted lines.
@@ 7 Simulations
In the previous sections, we have constructed $\gamma$- and $(\gamma,\gamma)$-MERGs, which provably achieve maximum $r$- and $(r,s)$-robustness with the fewest edges possible. In this section, we demonstrate the robustness of these graphs through two different sets of simulations (first in Sec. 7.1 and second in Sec. 7.2). We run each simulation from time step $t=0$ to $t=30$. In each simulation, the normals agents in a network run the W-MSR algorithm to perform consensus in the presence of $F$-total misbehaving agents, which transmit adversarial values to disrupt consensus. The simulation results are shown in Figure 5 and Figure 6, where the states of normal agents are plotted in solid colored lines, while states of misbehaving agents are plotted in red dotted lines.
@@ 7.1 Consensus Performance
In the first set of simulations, we demonstrate that in the presence of $F$ malicious agents, indexed by $i\in\{0,\dots,F-1\}$, normal agents $i\in\{F,\dots,n-1\}$ in $\gamma$- and $(\gamma,\gamma)$-MERGs are guaranteed to achieve resilient consensus. Each normal agent $i$’s initial state $x_{i}[0]\in\mathbb{R}$ is randomly generated on the interval $[-1000,1000]$. At time step $t\geq 0$, a malicious agent $i$ transmits an adversarial value: $1080\cos(t/5)$ if $i$ is even, and $1080\sin(t/5)$ if $i$ is odd.
We consider networks with $n=49$ and $n=50$ agents in all simulations. Maximum $r$-robustness any network with $49$ and $50$ agents can achieve is $r=\lceil\frac{n}{2}\rceil=25$-robustness. Thus, using Theorem 4.16, we construct $25$-MERGs with $49$ and $50$ nodes. Note these graphs achieve $25$-robustness with the fewest edges possible for a given number of nodes, and it can tolerate at most $F=12$ malicious agents. Thus, we run consensus simulations on these network graphs with $12$-total malicious agents, and the results are shown in Figure 5 (a) and (b). Similarly, the maximum $(r,s)$-robustness any network with $49$ and $50$ agents can achieve is $(25,25)$-robustness. Thus, we use Theorem 6.26 to construct $(25,25)$-MERGs with $49$ and $50$ agents. These graphs can tolerate up to $F=24$ malicious agents. We run consensus simulations on these network graphs with $24$-total malicious agents, and the results are shown in Figure 5 (c) and (d). As shown in Figure 5, both $\gamma$- and $(\gamma,\gamma)$-MERGs allow all normal agents to successfully achieve consensus within the convex hull of their initial values through the W-MSR, despite the presence of $F$ malicious agents.

### v2

@@ 6 Construction of $(\gamma,\gamma)$-MERGs
In the previous section, we presented the lower bounds on edge counts that any graph with $n$ nodes must satisfy in order to achieve $(\gamma,\gamma)$-robustness, where $\gamma=\lceil\frac{n}{2}\rceil$. Now, we present our second main result by constructing $(\gamma,\gamma)$-MERGs, or graphs that achieve $(\gamma,\gamma)$-robustness with a minimal set of edges:
Theorem 2.
@@
Proof.
Odd $n$: Because all $2\gamma-1$ nodes have $2(\gamma-1)$ neighbors, that means $\mathcal{G}$ forms a complete graph. Therefore by Lemma 6, $\mathcal{G}$ is a $(\gamma,\gamma)$-MERG when $n$ is odd.
Even $\mathbf{n}$: To prove that $\mathcal{G}$ is $(\gamma,\gamma)$-MERG for even values of $n$, we need to prove two things: (i) it is $(\gamma,\gamma)$-robust and (ii) its edge count $|\mathcal{E}|$ equals to the lower bounds specified in Lemma 8.
First, we prove its robustness. WLOG, we form disjoint non-empty subsets $\mathcal{S}_{1},\mathcal{S}_{2}\subset\mathcal{V}$ such that $|\mathcal{S}_{1}|\leq|\mathcal{S}_{2}|$, which implies $1\leq|\mathcal{S}_{1}|\leq\gamma$. By construction, there are at least $\gamma$ nodes in $\mathcal{G}$ that are connected to all other nodes (i.e., they has edges with $2\gamma-1$ nodes). There are two different cases:
@@
2) If $|\mathcal{S}_{1}|=|\mathcal{S}_{2}|=\gamma$, note that we have at least $\gamma$ nodes that have $2\gamma-1$ neighbors. Thus, however we form $\mathcal{S}_{1}$ and $\mathcal{S}_{2}$, there will be at least $\gamma=\frac{n}{2}$ nodes in $\mathcal{S}_{1}$ and $\mathcal{S}_{2}$ combined that have at least $2\gamma-1-(\gamma-1)=\gamma$ neighbors outside of $\mathcal{S}_{1}$ and $\mathcal{S}_{2}$, respectively. This satisfies the third condition given in Definition 8. Since $\mathcal{G}$ satisfies Definition 8 with $r=s=\gamma$ in all cases, $\mathcal{G}$ is $(\gamma,\gamma)$-robust.
Now we show that $\mathcal{E}$ is a minimal set. We have a total degree of $4\gamma(\gamma-1)+K$, where $K=\gamma+1$ if $\gamma$ is odd and or $K=\gamma$ if $\gamma$ is even. Thus, the total number of edges is $|\mathcal{E}|=\frac{4\gamma(\gamma-1)+K}{2}=\frac{K}{2}+2\gamma(\gamma-1)$, which equals the lower bound given in Lemma 8. ∎
Theorem 2 addresses Problem 2 by constructing $(\gamma,\gamma)$-MERGs, graphs that achieve maximum $(r,s)$-robustness with a minimal set of edges. Similar to $\gamma$-MERGs, these graphs also have as many edges as specified by the lower bounds in Lemmas 6 and 8. This shows that the lower bounds given in the lemmas are tight. In fact, removing any edge from $(\gamma,\gamma)$-MERGs immediately reduces the graphs’ robustness and thus resilience to misbehaving agents.
Examples of $(\gamma,\gamma)$-MERGs as the results of Theorem 2 are visualized in Figure 4. The graph on the left has $n=9$ nodes and forms a complete graph to achieve $(5,5)$-robustness, as discussed in Lemma 6. On the contrary, the graph on the right has $n=10$ nodes and is $(5,5)$-robust.
Furthermore, we highlight that the construction mechanisms given in Theorem 2 are necessary and sufficient in a sense that they encompass all the possible variations of $(\gamma,\gamma)$-MERGs.
@@
Let $\mathcal{G}=(\mathcal{V},\mathcal{E})$ be a graph of $n$ nodes, and let $\gamma=\lceil\frac{n}{2}\rceil$. Then, $\mathcal{G}$ is $(\gamma,\gamma)$-robust if and only if $\mathcal{G}$ contains a graph $\mathcal{G}^{\prime}=(\mathcal{V},\mathcal{E}^{\prime})$ from Theorem 2 as a spanning subgraph (i.e., $\mathcal{E}^{\prime}\subseteq\mathcal{E}$).
Proof.
@@
(necessary) Odd $n$: It is shown in Lemma 6 that complete graph is the only graph with $n$ nodes that is $(\gamma,s)$-robust where $s\geq\lfloor\frac{n}{2}\rfloor$. This indicates that $\mathcal{G}$ must be a complete graph, which means $\mathcal{E}^{\prime}=\mathcal{E}$, as $\mathcal{G}^{\prime}$ from Theorem 2 is also a complete graph.
@@
Even $n$: Here we prove the claim by contradiction. Assume by contradiction that $\mathcal{G}$ is $(\gamma,\gamma)$-robust and $\mathcal{E}^{\prime}\not\subseteq\mathcal{E}$. Let $\mathcal{N}_{i}$ and $\mathcal{N}_{i}^{\prime}$ be the neighbor sets of agent $i\in\mathcal{V}$ in $\mathcal{G}$ and $\mathcal{G}^{\prime}$, respectively. Then, we define $\mathcal{M}=\{i\in\mathcal{V}\mid|\mathcal{N}_{i}|=2\gamma-1\}$, $\mathcal{P}=\{i\in\mathcal{V}\mid|\mathcal{N}_{i}|=2\gamma-2\}$, $\mathcal{M}^{\prime}=\{i\in\mathcal{V}\mid|\mathcal{N}_{i}^{\prime}|=2\gamma-1\}$, and $\mathcal{P}^{\prime}=\{i\in\mathcal{V}\mid|\mathcal{N}_{i}^{\prime}|=2\gamma-2\}$.
@@
Note $\mathcal{G}^{\prime}$ has $|\mathcal{M}^{\prime}|=2\cdot\lceil\frac{\gamma}{2}\rceil$ nodes with degree of $2\gamma-1$ and $|\mathcal{P}^{\prime}|=2\gamma-|\mathcal{M}|$ nodes with degree of $2\gamma-2$. By Corollary 2, all nodes cannot have less than $2\gamma-2$ neighbors for $\mathcal{G}$ to be $(\gamma,\gamma)$-robust. Furthermore, they cannot have more than $2\gamma-1$ neighbors. What that means is that the only way for $\mathcal{E}^{\prime}\not\subseteq\mathcal{E}$ is to have $|\mathcal{M}|<|\mathcal{M}^{\prime}|$. Then, $|\mathcal{E}|<2\gamma(\gamma-1)+\lceil\frac{\gamma}{2}\rceil$. However, that is a contradiction, as for $\mathcal{G}$ to be $(\gamma,\gamma)$-robust, $|\mathcal{E}|\geq 2\gamma(\gamma-1)+\lceil\frac{\gamma}{2}\rceil$ by Lemma 8.
@@
(sufficiency) We know that $\mathcal{G}^{\prime}$ is a $(\gamma,\gamma)$-MERG, i.e., $(\gamma,\gamma)$-robust graph with a minimal set of edges. Since $(r,s)$-robustness is monotonic with respect to number of edges [4, Lemma 3], if $\mathcal{E}^{\prime}\subseteq\mathcal{E}$, $\mathcal{G}$ must be $(\gamma,\gamma)$-robust. ∎
It follows from 1 that every $(\gamma,\gamma)$-MERG - that is, every $(\gamma,\gamma)$-robust graph with a minimal edge set - can be constructed using the procedure in Theorem 2. Consequently, it is possible to use 1 to quickly verify whether a given graph $\mathcal{G}$ is $(\gamma,\gamma)$-robust by checking whether it contains $(\gamma,\gamma)$-MERGs from Theorem 2 as a spanning subgraph. While finding such subgraphs is in general NP-complete, many efficient algorithms exist (e.g., [53, 54]).
@@
Figure 6: Simulation results for the second set of simulations. Consensus with $F=2$ Byzantine agents is performed on 55-MERGs with $n=9$ (panels (a) and (e)) and $n=10$ (panels (b) and (f)), both without and with an edge removal, respectively. Similarly, consensus with $F=4$ malicious agents is executed on $(5,5)$-MERGs with $n=9$ (panels (c) and (g)) and $n=10$ (panels (d) and (h)), also without and with an edge removal, respectively. The states of normal agents are shown as solid lines, while those of misbehaving agents are omitted for visual clarity.
7 Comparison with Other Graphs
Table 1: Number of edges required to achieve maximum robustness for different graph construction methods as the number of nodes $n$ increases.
$n$ Graphs Edge Count Edge Count
($\gamma$-robustness) ($(\gamma,\gamma)$-robustness)
5 $k$-Circulant [49] 10 10
$F$-Elemental [37] 9 -
Ours 9 10
30 $k$-Circulant [49] 435 435
$F$-Elemental [37] - -
Ours 323 428
105 $k$-Circulant [49] 5460 5460
$F$-Elemental [37] 4134 -
Ours 4134 5460
300 $k$-Circulant [49] 44850 44850
$F$-Elemental [37] - -
Ours 33601 44775
In this section, we compare $\gamma$- and $(\gamma,\gamma)$-MERGs from Theorems 1 and 2 with other existing robust graphs in terms of edge counts. The goal is to highlight that our construction methods require less edges than other existing methods. We focus on two classes of graphs for comparison: undirected $k$-circulant graphs [49, 8] and $F$-elemental graphs [37]. Table 1 summarizes the number of edges required to achieve maximum robustness for each class of graphs as the number of nodes $n$ increases.
Undirected $k$-circulant graph is $\lceil k/2\rceil$-robust and at least $(\lfloor(k+2)/2\rfloor,\lfloor(k+2)/2\rfloor)$-robust for even $k$ and $(\lfloor(k+1)/2\rfloor,\lfloor(k+1)/2\rfloor)$-robust for odd $k$ [49, 55]. Consequently, achieving the maximum robustness level requires the graph to be fully connected, leading to a high number of edges, as shown in Table 1. On the other hand, $F$-Elemental graphs constructed with [37, Prop. 1] (discussed in Remark 3) tend to use as few edges as ours. However, these graphs are defined only $r$-robust graphs with odd $n$ and provide no analysis of edge minimality. In contrast, as supported by our analyses in Sections 3 and 5, our graphs constructed according to Theorems 1 and 2 consistently use the fewest edges for any $n$ and for both $r$- and $(r,s)$-robustness.
@@ 8 Simulations
In the previous sections, we have established systematic ways to construct $\gamma$- and $(\gamma,\gamma)$-MERGs, which provably achieve maximum $r$- and $(r,s)$-robustness with the fewest edges possible. We now validate the correctness of our results through two sets of simulations (presented in Section 8.1 and Section 8.2). While $\gamma$- and $(\gamma,\gamma)$-MERGs support a variety of resilient algorithms, including estimation, optimization, and learning [15, 16, 18] as well as different resilient consensus algorithms [43, 39, 56, 57], here we focus specifically on resilient consensus via the W-MSR algorithm.
For all simulations, normal agents in the network execute the W-MSR algorithm to achieve consensus despite the presence of misbehaving agents, which transmit adversarial values to disrupt consensus. Through the W-MSR algorithm, normal agents’ states are guaranteed to converge to a common value within the convex hull of their initial conditions, irrespective of the behaviors of the misbehaving agents [4]. The results are shown in Figures 5 and 6.
@@ 8.1 Consensus Performance
In the first set of simulations, we demonstrate that in the presence of $F$ malicious agents, indexed by $i\in\{0,\dots,F-1\}$, normal agents $i\in\{F,\dots,n-1\}$ in $\gamma$- and $(\gamma,\gamma)$-MERGs are guaranteed to achieve resilient consensus. Each normal agent $i$’s initial state $x_{i}[0]\in\mathbb{R}$ is randomly generated on the interval [1000,1000][-1000,1000]. At time step $t\geq 0$, a malicious agent $i$ transmits an adversarial value: $1080\cos(t/5)$ if $i$ is even, and $1080\sin(t/5)$ if $i$ is odd.
We consider networks with $n=49$ and $n=50$ agents in all simulations. Maximum $r$-robustness any network with $49$ and $50$ agents achieve is $r=\lceil\frac{n}{2}\rceil=25$-robustness. Thus, using Theorem 1, we construct $25$-MERGs with $49$ and $50$ nodes. Note these graphs achieve $25$-robustness with the fewest edges possible for a given number of nodes, and it tolerates at most $F=12$-local malicious agents. Thus, we run consensus simulations on these network graphs with $12$ malicious agents, and the results are shown in Figure 5 (a) and (b). Similarly, the maximum possible $(r,s)$-robustness any network with $49$ and $50$ agents achieves is $(25,25)$-robustness. Thus, we use Theorem 2 to construct $(25,25)$-MERGs with $49$ and $50$ agents. These graphs tolerate up to $F=24$-total malicious agents. We run consensus simulations on these network graphs with $24$ malicious agents, and the results are shown in Figure 5 (c) and (d). As shown in Figure 5, both $\gamma$- and $(\gamma,\gamma)$-MERGs allow all normal agents to successfully achieve consensus within the convex hull of their initial values through the W-MSR algorithm with $F$ malicious agents.